OpenStudy (anonymous):

EVALUATE USING ALGEBRA LIM X^2 +2 / X -1 AS X GOES TO NEGATIVE INFINITY

OpenStudy (anonymous):

The strategy for solving this is the same as the one I solved before, except the square root is gone and the limit boundary is negative infinity which will change the answer. \[\lim_{x \rightarrow -\infty}(x^{2}+2)/(x-1)\] (factor out x^2 in the numerator and x in denominator) \[\lim_{x \rightarrow -\infty}x^{2}(1+2/x^{2})/x(1-1/x) \] \[\lim_{x \rightarrow -\infty}x(1+2/x^{2})/(1-1/x)\] \[\lim_{x \rightarrow -\infty}x(1+0)/(1+0) \] \[\lim_{x \rightarrow -\infty}x=-\infty\] So the limit is \[-\infty\]